OpenStudy (anonymous):

23. Given , explain how to construct a square with sides of length AB. 24. Is the line through points P(–8, –10) and Q(–5, –12) perpendicular to the line through points R(9, –6) and S(17, –5)? Explain.

OpenStudy (anonymous):

#1) I need more detail. #2) Find the slope of both lines. Of they are opposite reciprocals of each other, then they are perpendicular.

OpenStudy (anonymous):

\[Slope = \frac{ y2 - y1 }{ x2 - x1 } \]

OpenStudy (anonymous):

The slope for line PQ is:\[PQ \to Slope = \frac{ -12 - (-10) }{ -5 - (-8) } = \frac{ -12 + 10 }{ -5 + 8 } = \frac{ 2 }{ 3 }\] Find the slope for line RS, if its -3/2 the lines are perpendicular.

OpenStudy (anonymous):

Does this help @sydneyyy ?

OpenStudy (anonymous):

Need help on this if you can